# Do Determinants of EV Purchase Intent vary across the Spectrum? Evidence from Bayesian Analysis of US Survey Data

### v2

@@ 4.2 Results from Ordinal Quantile Models
We now present the results from ordinal quantile models (Rahman, 2016) that shift attention from average probabilities of outcomes to those at specific quantiles, conditional on the covariates. To our knowledge, this aspect has not been examined in prior work on EVs. Table 4 presents the posterior means and standard deviations of the parameters from the ordinal quantile models estimated separately for each survey; along with the logarithm of marginal likelihood which can be used for Bayesian model comparison (Maheshwari and Rahman, 2023). Note that the coefficients in Table 4 are not directly comparable to those in Table 2 for two main reasons: (a) the models operate on different scales due to differences in the variance of the error terms, and (b) they address distinct inferential goals-the ordinal probit model focuses on the average probability of outcomes, whereas the ordinal quantile model targets the probability of outcomes at specific quantiles.
@@ 5 Discussion
Our findings demonstrate how targeted strategies may be able to accelerate EV adoption. Exposure to EV information is associated with a 10.2 percentage point higher probability of reporting strong intent, with larger differences among respondents at the upper end of the intent distribution. It is also associated with an 8.5 point lower probability of reporting no intent. However, only 37.7% of respondents report high exposure, pointing to an opportunity for broader outreach through media, industry, and education. The smaller covariate effect of prior exposure at the 20th percentile suggests that simply raising awareness may not persuade consumers with low purchase interest. In contrast, for those already inclined toward adoption, providing practical, consumer-focused information—such as total cost of ownership, available incentives, and charging logistics—may influence purchase decisions.
Belief that EVs are better for the environment than gas-powered vehicles is associated with a 15.5 percentage point higher probability of reporting strong intent, with larger differences at the upper end of the distribution. It is also associated with a 19.5 point lower probability of reporting no interest, with larger differences at the lower end of the distribution. This pattern suggests that belief in EVs’ environmental benefits reinforces commitment among likely adopters and reduces skepticism among consumers with lower purchase interest. In the 2021 survey, 71.8% of respondents expressed this belief. Emphasizing environmental benefits—such as improved air quality—may resonate with both likely adopters and more skeptical respondents, broadening support for EV adoption.
Confidence in charging infrastructure is one of the strongest predictors of purchase intent. Respondents who express confidence that the US will build sufficient infrastructure are 19.1 percentage points more likely to report strong intent and 18.8 points less likely to report no interest. Quantile estimates suggest that confidence both strengthens commitment among those already inclined toward adoption and reduces outright rejection among those with lower purchase intent. Despite the importance of this factor, only 17.5% of respondents in the 2023 survey reported high confidence. Policymakers and industry can address this gap through continued investment in charging infrastructure-such as the National Electric Vehicle Infrastructure (NEVI) program-to improve affordability, accessibility, and reliability, as well as through effective communication of these efforts and of a long-term vision for the charging network.
Belief that the government is doing too much to address climate change is associated with a 14.8–19.7 point lower probability of strong intent and a 19.7–33.4 point higher probability of no interest. In the 2023 survey, 24.1% of respondents held this view. These findings suggest that efforts to promote adoption can be more effective if they emphasize benefits that resonate across ideological lines. Rather than positioning EVs primarily as a climate solution, highlighting other benefits—such as long-term cost savings and air quality improvements—may engage a broader range of potential buyers. Our finding that Democrats are more likely to express stronger interest in EVs, further highlights the role of political identity and the value of emphasizing non-partisan benefits to broaden appeal. Moreover, for consumers wary of climate-focused policies, endorsements from non-governmental entities, such as automakers, dealerships, and industry groups, may be more persuasive than government-led campaigns.
The results for other covariates point to further implications. Prior EV ownership is the strongest predictor of purchase intent, underscoring the importance of reducing barriers for first-time buyers. Potential options include subsidies, test-drive programs, and community-based education. Younger, urban, educated, and higher-income individuals report greater interest in EVs, suggesting that targeting information or advertising to such individuals may help raise adoption rates. By contrast, lower interest among women, US-born, and Black respondents may reflect cultural, structural, or informational barriers that could be addressed through inclusive policy design and outreach.
@@ 6 Conclusion
EVs are increasingly regarded as a cornerstone of transportation decarbonization and broader climate change mitigation efforts (Bhat et al., 2025). However, federal support has varied along partisan lines, with Democratic administration promoting EV adoption and Republican administration prioritizing deregulation and rolling back environmental standards. Amid these policy shifts, consumer interest remains a key driver of both market trends and national policy. Although public opinion is known to influence US policymaking (Burstein, 2003; Shapiro, 2011), yet no prior research has analyzed EV purchase sentiment using a nationally representative sample over multiple years or quantified the impact of covariates. In addition, no studies have applied quantile analysis to examine how different factors influence the distribution of intention to purchase EVs.
This study addresses these gaps by analyzing public attitudes toward EVs using three nationally representative Pew Research Center surveys from 2021-2023. We take the Bayesian route and model EV purchase intent with ordinal probit and ordinal quantile regression, incorporating novel predictors, prior EV ownership, political affiliation, and demographic factors. Results from the ordinal probit models show that strong interest in purchasing an EV is associated with greater exposure to EV information, belief in their environmental benefits, prior EV or hybrid ownership, and confidence in charging infrastructure. Additional positive predictors include being under 50, higher education and income, metropolitan residence, living in the West, and Democratic Party affiliation. Conversely, lower interest is linked to skepticism about government climate efforts, identifying as female or Black, living in the Midwest, and US-born status.
Ordinal quantile analysis reveals that the effects of key factors on EV purchase intent vary across the distribution. Respondents who are knowledgeable about EVs, view them as environmentally beneficial, have confidence in charging infrastructure, own an EV, or identify as Democrats are more likely to express strong interest at higher quantiles of EV purchase intent. These same factors are associated with lower likelihood of no interest at lower quantiles. Conversely, those who believe the government is overreaching on climate policy are more likely to express no interest at lower quantiles and less likely to show strong interest at higher ones. These findings highlight how covariate effects differ across levels of purchase intent, thereby providing a more nuanced understanding than the ordinal probit model.
The findings offer valuable guidance for policymakers, industry stakeholders, and advertisers seeking to bolster EV adoption. Confidence in charging infrastructure is one of the strongest predictors of purchase intent—associated with greater interest among likely adopters and reduced rejection among skeptics—yet less than 20% respondents report high confidence. Infrastructure investment, along with clearer communication about availability, reliability and expansion plans, may help close this gap. Exposure to EV information is also linked to stronger intent and reduced disinterest, though only 37.7% of respondents report high exposure—highlighting the value of broader outreach through media, industry, and education. Belief that EVs are environmentally superior to gas-powered vehicles is similarly associated with higher intent and lower disinterest, with effects varying across the intent distribution; messaging focused on benefits such as improved air quality may therefore be effective. Perceptions that the government is doing too much on climate change are negatively associated with intent, suggesting that outreach emphasizing cost, convenience, or energy security—particularly from non-government sources—may have broader appeal.

### v3

@@ 4.2 Results from Ordinal Quantile Models
We now present the results from ordinal quantile models [46] that shift attention from average probabilities of outcomes to those at specific quantiles, conditional on the covariates. To our knowledge, this aspect has not been examined in prior work on EVs. Table 4 presents the posterior means and standard deviations of the parameters from the ordinal quantile models estimated separately for each survey; along with the logarithm of marginal likelihood which can be used for Bayesian model comparison [41]. Note that the coefficients in Table 4 are not directly comparable to those in Table 2 for two main reasons: (a) the models operate on different scales due to differences in the variance of the error terms, and (b) they address distinct inferential goals-the ordinal probit model focuses on the average probability of outcomes, whereas the ordinal quantile model targets the probability of outcomes at specific quantiles.
@@ 5 Discussion
Our findings point to several targeted strategies that may accelerate EV adoption by clarifying how beliefs and expectations shape purchase intent. First, exposure to EV-related information is associated with approximately a 10 percentage point higher purchase intent. This pattern is notable given that fewer than 40 percent of respondents report high exposure, suggesting scope to expand the reach of EV-related information among prospective buyers. At the same time, smaller estimated effects at the lower end of the intent distribution indicate that increased awareness alone is unlikely to materially increase purchase intent among consumers with low baseline interest. Together, these results suggest that information provision may be most effective when targeted toward consumers already considering an EV. Such information could focus on practical considerations such as costs, incentives, and charging logistics. These results complement prior work that infers information exposure through peers or media [49, 58] by relying on respondents’ self-reported exposure and by highlighting heterogeneity across the intent distribution rather than average effects.
Second, belief that EVs have lower environmental impacts than gasoline vehicles is associated with higher purchase intent and lower reported disinterest. This result complements prior work linking pro-environmental attitudes to EV support [e.g., 22, 42] by showing that adoption decisions are shaped not only by general environmental orientation, but also by consumers’ perceptions of EVs’ environmental benefits relative to gasoline vehicles. It also extends evidence from smaller-sample studies examining the effects of perceived environmental benefits relative to gasoline vehicles to a larger, nationally representative US sample [36, 29]. Survey responses indicate that a large majority of respondents believe EVs are environmentally preferable, but do not capture beliefs about the magnitude of those benefits. This measurement limitation prevents us from assessing how much additional adoption could be encouraged by shifting beliefs, and highlights a need for future research on how consumers perceive the size of environmental benefits and how those perceptions respond to information.
Third, confidence that sufficient charging infrastructure will be built is strongly associated with purchase intent, yet fewer than 20 percent of respondents report high confidence. Against the backdrop of prior work emphasizing physical access to charging infrastructure [10, 55], our results point to expectations about the future charging network as a distinct channel shaping adoption decisions. These findings underscore the role of policy stability and credible long-term commitments in shaping expectations about the development of the charging network. Concerns about charger reliability and maintenance may also weaken confidence in future access even as new infrastructure is deployed. Such concerns were salient during the period covered by our survey data, and industry surveys reported that a substantial share of public-charging visits did not result in a successful charge due to charger unavailability or inoperability [24]. Policymakers and industry can help address this confidence gap through continued investment in charging infrastructure—such as through the National Electric Vehicle Infrastructure (NEVI) program—together with clear communication of deployment and maintenance plans.
Fourth, belief that the government is doing too much to address climate change is associated with lower EV purchase intent and is held by a non-trivial share of respondents. Our results complement prior findings linking EV interest to political affiliation [42, 53] by indicating that evaluations of climate policy itself, in addition to partisan affiliation, shape EV adoption decisions. This pattern implies that framing EVs primarily as a climate policy instrument is unlikely to resonate with consumers who are skeptical of climate-focused policies. For these consumers, information about other EV attributes—such as operating costs, performance, or local air quality—that may be less salient ex ante can still influence adoption decisions. Moreover, information conveyed by non-governmental actors may carry greater credibility than government-led messaging.
Finally, prior EV ownership is strongly associated with stated purchase intent, consistent with learning or reduced uncertainty following initial adoption. This pattern suggests that policies lowering barriers for first-time buyers may influence adoption dynamics beyond the initial purchase. At the same time, systematic differences in stated interest across demographic groups may point to disparities in exposure, access, or information, highlighting the importance of inclusive policy design and outreach.
@@ 6 Conclusion
Electric vehicles (EVs) are widely viewed as a cornerstone of transportation decarbonization and climate change mitigation, yet federal support for EV adoption in the US has fluctuated sharply across administrations, reflecting broader partisan divides. In this context, consumer purchase sentiment plays a critical role in shaping both market outcomes and the political feasibility of sustained EV policy. Although public opinion is known to influence US policy making, existing research has not examined EV purchase intent using nationally representative data over multiple years, nor has it quantified how certain determinants operate across different levels of purchase interest. This study addresses these gaps by analyzing three nationally representative Pew Research Center surveys from 2021-2023 using Bayesian ordinal probit and ordinal quantile regression models.
The empirical results indicate that EV purchase intent is positively associated with greater exposure to EV-related information, stronger beliefs in the environmental benefits of EVs, and confidence in the development of charging infrastructure. Political affiliation and demographic characteristics also play a significant role, with higher intent observed among younger, more educated, and higher-income individuals, metropolitan residents, Democrats, and those living in the western United States. In contrast, lower purchase intent is associated with skepticism toward government climate action, as well as with certain demographic characteristics, underscoring persistent heterogeneity in attitudes toward EV adoption across population subgroups.
The ordinal quantile analysis further reveals substantial heterogeneity in the determinants of EV purchase intent across the distribution. Greater EV knowledge, perceptions of environmental benefits, confidence in charging infrastructure, prior EV ownership, and Democratic Party affiliation increase the likelihood of strong purchase interest at higher quantiles while simultaneously reducing outright rejection at lower quantiles, whereas perceptions that the government is overreaching on climate policy are associated with lower interest throughout the distribution. These results provide a more nuanced perspective than mean-based ordinal models and yield clear policy and market insights: confidence in charging infrastructure emerges as one of the strongest predictors of adoption despite being reported by fewer than 20% of respondents, underscoring the importance of infrastructure investment and transparent communication about availability and reliability. Similarly, broader exposure to EV-related information and messaging emphasizing environmental benefits—such as improved air quality—are associated with higher intent and reduced disinterest, while outreach that stresses cost, convenience, or energy security, particularly from non-government sources, may be more effective among climate-policy skeptics.